A display of cans on a grocery shelf consists of 20 cans on the bottom, 18 cans in the next row, and so on in an arithmetic sequence, until the top row has 4 cans. How many cans, in total, are in the display?
please use the equation needed so i know how to do it.
thank you:)

Answer by stanbon(75887)

A display of cans on a grocery shelf consists of 20 cans on the bottom, 18 cans in the next row, and so on in an arithmetic sequence, until the top row has 4 cans. How many cans, in total, are in the display?
-----------------------------
20 + 18 + ... + 6 + 4
---------------------------
a = 20 ; d = -2 ; n = ?
--------
4 = 20 + (n-1)(-2)
-16 = (n-1)(-2)
n-1 = 8
n = 9
-------
S(9) = (9/2)(20+4)
S(9) = 4.5*24
S(9) = 108
===================
cheers,
Stan H.